Using the financial statements for the Snider Corporation, calculate the 13 basic ratios found in the chapter SNIDER CORPORATION Balance Sheet December 31, 20X1 Assets Current assets: Cash Marketable securities Accounts receivable (net) Inventory $ 54,200 28,800 180,000 244,000 $507,000 60,400 Total current assets Investments $658,000 255,000 Plant and equipment. Less: Accumulated depreciation Net plant and equipment 403,000 $970,400 Total assets Liabilities and Stockholders Equity Current liabilities: Accounts payable Notes payable Accrued taxes $94,800 78,400 14,500 $187,700 Total current liabilities Long-term liabilities: Bonds payable Total liabilities 159,800 $347,500 Stockholders equity Preferred stock, $50 par value Common stock, $1 par value Capital paid in excess of par Retained earnings $100,000 80,000 190,000 252,900 $622,900 Total stockholders equity Total liabilties and stockholders equity $970,400 SNIDER CORPORATION Income Statement For the Year Ending December 31, 20X1 Sales (on credit) Cost of goods sold Gross proft Selling and administrative expenses Operating profit (EBIT) Interest expense Earnings before taxes (EBT) Taxes $2,070,000 1.377,000 $ 693,000 505,000* $ 188,000 34,200 153,800 85,800 $ 68,000 Earnings after taxes (EAT) Includes $35,800 in lease payments Using the above financial statements for the Snider Corporation, calculate the following ratios. a. Profitability ratios. (Do not round intermediate calculations. Input your answers as a percent rounded to 2 decimal places.) Profitability Ratios Profit margin Return on assets investment Return on equitb. Assets utilization ratios. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Assets Utilization Ratios Receivable turnover Average collection Inventory turnover Fixed asset turnover Total asset imes days imes imes c. Liquidity ratios. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Liquidity Ratios Current ratio Quick ratio times imes d. Debt utilization ratios. (Do not round intermediate calculations. Input your debt to total assets answer as a percent rounded to 2 decimal places. Round your other answers to 2 decimal places.) Debt Utilization Ratios Debt to total assets Times interest earned Fixed charge covera imes times
Answer #1

a)
Profit margin=net income/sales
=68000/2070000=3.29%
Return on assets= net income/assets
=68000/970400=7.01%
Return on equity= Net income/equity
=68000/622900=10.92%

b)
Receivable Turnover= sales/account receivable
=2070000/180000=11.50
Average collection period=365/Receivable Turnover
=365/11.5=31,74
Inventory turnover=cost of goods sold/inventory
=1377000/244000=5.64
Fixed asset turnover= sales/fixed assets
=2070000/403000=5.14
total asset turnover= sales/total assets
=2070000/970400=2.13

c)
current ratio=current assets/current liabilites
=507000/187700=2.70
quick ratio=(current asstes-inventory)/current liabilites
=(507000-244000)/187700=1.40

d) Debt to total assets=debt/total assets
=347500/970400=35.81%
times interest earned=EBIT/interest
=188000/34200=5.50
Fixed charge coverage=(operating profit+lease expense)/(interest expense+lease expense)
=(188000+35800)/(34200+35800)=3.20